OpenStudy (anonymous):

pythagorean theorem a=7 b=4 c=?

OpenStudy (anonymous):

you can do this. do you know what 7 squared is?

OpenStudy (anonymous):

c=root(a^2+b^2) =root(49+16) =root(65)

OpenStudy (anonymous):

\[\sqrt{65}\] = 8.0622 dhashni is right...